A circular inch and square inch are used in diameter and area of electrical wiring.
A circular inch is a circle 1 inch in diameter.
1 circular inch = .7854 sq. in.: Pi * r^2; 3.1416 * .5 * .5 = .7854 sq. in.
How do I calculate 1 sq. in. = x circular in. ?

Answer by greenestamps(13206)



The same way you perform more familiar conversions between units....

Example 1: convert 3 hours to minutes:
              60 minutes
   3 hours * ------------ = 180 minutes
                1 hour

Example 2: convert 440 yards to miles:
                  1 mile
   440 yards * ----------- = 440/1760 miles = 1/4 mile
               1760 yards

Your example: convert 1 square inch to circular inches:
                       1 circular inch
    1 square inch * ---------------------- = 1/0.7854 circular inches
                     0.7854 square inches

Use a calculator to get approximately 1.2732.